Question 2 a. What is GDP? Explain the three methods that the Australian Bureau of Statistics uses to calculate GDP.

b. If in 2016 the CPI is 100 and nominal wages are $500 and in 2017 the CPI is 120 and nominal wages are $550. What is the level of price inflation from 2016 to 2017? Explain whether real wages have increased from 2016 to 2017?

c. Define structural, frictional and cyclical unemployment. Which of these types of unemployment do the 'long-term unemployed' belong?

d. Explain why the growth in real per capita GDP is a more appropriate measure of economic growth that nominal GDP for the whole country. e. In an 'economic boom' what is likely to happen to inflation, unemployment and the participation rate? Explain why.
Answer #1
  1. GDP is the gross domestic product. It means that total production that happens within the territory of the country.

Three methods of calculating GDP:

  1. Income approach
  2. Expenditure approach
  3. Production approach

b). Price inflation: it is the increase in the price over a period of time.

= 550-500/120-100 = 50/20 = 2.5

The real wages have increased over e period of 1 year.

c). Structural unemployment:

This happens due to structural changes in dynamic economy making some workers go out of job. These changes are a continuous process.

Example: sectoral changes in economy like in change in technology.

Frictional unemployment:

Frictional unemployment is the condition where some firms close down leads to the reduction in the employment.

Cyclical unemployment:

It is related to the cyclical trend and business cycle. When the business cycle is at peak then the unemployment would be less and vice versa.

Structural unemployment is the 'long-term unemployed'.

D). real GDP shows the more appropriate measure because real GDP shows the increase in production while the nominal GDP changes due to production price.

e). in the boom conditions, the unemployment would be low, inflation lead to boom condition. Boom also increases the participation rate.